# The Great Ban: Efficacy and Unintended Consequences of a Massive Deplatforming Operation on Reddit

### v1

@@ Introduction
Table 2: Subreddit-wise median toxicity scores for users who abandoned Reddit after the ban (ABA) and for those who remained. For the latter, toxicity scores are computed both before (BEF) and after (AFT) the ban. The ABA vs BEF column shows that users who abandoned the platform were more toxic than those who remained. The BEF vs AFT column shows that users who remained active experienced a modest toxicity reduction after the ban.
RQ1: Effectiveness of The Great Ban
@@
Abandoning users. Table 1 highlights a difference of 2,577 users (15.6%) between the OUT-BEFORE and OUT-AFTER datasets, corresponding to users who became inactive on Reddit after The Great Ban. Given that such users were consistently active before the ban, but did not post a single comment in the 7 months following it, we conclude that they abandoned Reddit, possibly migrating to other platforms (Horta Ribeiro et al. 2021). This result thus highlights a first straightforward effect of The Great Ban. In order to draw more insights into this finding, we compare the pre-ban subreddit-wise toxicity of the abandoning users with that of the users who remained active on the platform. Table 2 shows that in 14 out of 15 subreddits, users who later abandoned the platform were more toxic than those who remained after the ban. In other words, toxic users were more likely to abandon the platform after the intervention than less toxic ones. For 12 subreddits the difference in toxicity between abandoning and remaining users is statistically significant ($p<0.05$), according to a non-parametric Mann-Whitney test for unpaired data. Moreover, in 12 subreddits, abandoning users have larger mean absolute deviation (MAD) scores than remaining users, indicating greater variability in toxicity among the former user group.
Figure 3: User-level toxicity change after The Great Ban, for each active user. The slope chart in the central panel reveals a majority of red-colored rising lines, corresponding to a large number of users who drastically increased their toxicity. The beeswarm plot in the bottom panel confirms this effect, presenting more users in the right red-colored tail of the distribution than in the left blue-colored one. Boxplots present marginal distributions for abandoning and remaining user, before and after the intervention.
@@
Remaining users. In addition to causing some users to abandon the platform, the ban might also have caused toxicity changes in those users who remained. Table 2 reports subreddit-wise toxicity scores for the matched set of remaining users, before and after the ban. As shown, the remaining users were on average less toxic after the ban. Specifically, when aggregating results for each subreddit, users from all 15 subreddits slightly decreased their toxicity. The decrease is statistically significant for 11 subreddits ($p<0.1$), according to a non-parametric Wilcoxon test for paired data. Notably, effect sizes for the comparison between remaining users before vs after the ban are smaller than those of the comparison between abandoning vs remaining users. In fact, after The Great Ban, the overall toxicity decreased by 6.57% on average – a modest amount. Moreover, for 13 out of 15 subreddits, toxicity MAD values are larger after the ban than before. This result suggests that the ban increased the degree of variability in the behavior of the remaining users, which we further investigate in the following.
@@
User-level effects. So far, we provided platform- and community-level results, finding that The Great Ban caused 15.6% of core users to abandon the platform. Furthermore, those who remained exhibited a modest average toxicity reduction of 6.57%. To provide a thorough assessment of the intervention, here we also investigate user-level effects by computing the toxicity changes experienced by each of the 13,963 users who remained active on Reddit after the ban. The central panel of Figure 3 presents a slope chart of all user-level toxicity changes, independently of the subreddit in which users participated. Each line corresponds to a single user. Line slopes encode the amount of toxicity reduction or increment. Rising lines are colored with different shades of red depending on their slope, and denote users who increased their toxicity after the ban. Contrarily, decreasing lines are blue-colored and denote users who decreased their toxicity. Figure 3 also includes marginal boxplot toxicity distributions for remaining users before (left-hand side of the slope chart) and after (right-hand side) the ban. The leftmost boxplot presents the toxicity distribution for users who abandoned the platform. Finally, the bottom panel shows the distribution of user-level toxicity changes as a beeswarm plot. This latter plot is useful for highlighting outliers and for studying the two tails of the distribution – i.e., those related to marked toxicity increases (red dots, right-hand side of the beeswarm plot) and decreases (blue dots, left-hand side).
@@
The three boxplots of Figure 3 confirm the overall toxicity trends observed in Table 2. Abandoning users have the largest median toxicity, followed by remaining users before the ban. Out of the three user groups, the remaining users after the ban have the lowest median toxicity. The slope chart in Figure 3 depicts a majority of lines with relatively small positive or negative slopes. These correspond to users who exhibited small toxicity changes – either increases or decreases – after the ban. At the same time, however, the slope chart also features a remarkable number of steep lines, which are related to users who exhibited large toxicity changes. Specifically, we note an overwhelming majority of red-colored steep lines. This implies that, among users who exhibited large toxicity changes, the vast majority increased their toxicity. This result qualitatively describes an undesired effect of The Great Ban, which caused a non-negligible minority of users to become resentful or vindictive, thus exhibiting much more toxic behaviors. The points where the lines of the slope chart intersect the y axis on the right-hand side of the plot represent the toxicity of the users after the ban. The distribution of the post-ban toxicity scores is also depicted in the right-hand side boxplot. As shown in Figure 3 and as anticipated from the MAD values in Table 2, there is more variability in user toxicity after the ban. Figure 3 clarifies that this increased variability is caused by the increased toxicity of the resentful users.
@@
Figure 4: User-level toxicity changes for participants in r/the_donald (top) and r/consumeproduct (bottom).
RQ2: Extreme user reactions to The Great Ban
@@
Figure 5: Subreddit-wise changes in toxicity obtained by summing all individual decreasing and increasing contributions, for all users (left panel) and outlier users (right panel). Outlier users are those whose change in toxicity exceeds the threshold $k=0.25$. Based on this criterion, we found no outliers for r/hatecrimehoaxes, which is thus omitted from the right panel figure.
@@
Figure 6: Contribution of the outlier users to the increase/decrease of toxicity in each subreddit. Outlier users are defined as those whose individual change in toxicity exceeds the threshold $k$. As shown, in 12 out of 15 subreddits the outlier users caused large toxicity increases.
Figure 7: Relationship between user participation in the 15 banned subreddits and change in toxicity. The vast majority of users only participated in one or two subreddits. Moreover, participation in more/less subreddits is unrelated to changes in toxicity.
@@
In order to specifically investigate the behavior of outlier users – those who experienced marked changes in toxicity – we recompute the previous summations by only considering users whose change in toxicity exceeds a given threshold $k$: $|\Delta t(i)|>k$. This allows us to focus on those users who experienced extreme behavioral changes, be them increases or decreases in toxicity. For example, the right panel of Figure 5 repeats the previous analysis by applying the threshold $k=0.25$. As shown, we obtain strikingly different results. When only considering outlier users, in 13 out of 15 subreddits the contributions of those who increased or decreased their toxicity to the overall subreddit toxicity are very imbalanced. The only exceptions are r/gendercritical who presents relatively balanced contributions (55% vs 45%) and r/hatecrimehoaxes for which no user experienced a change in toxicity exceeding the threshold $k=0.25$. Moreover, in 11 subreddits toxicity increases greatly outweigh decreases. To generalize these findings, we repeated the analysis by varying the value of $k$ from 0 to 1, with a step of 0.01. Results are shown in Figure 6 where each line corresponds to a subreddit and reports the % difference $\Delta t^{+}-\Delta t^{-}$ between the contributions of the two tails, for each value of $k$. In the figure, only a few notable subreddits are highlighted, while the rest are grey-colored so as to reveal the overall trend. As shown, even for small values of the threshold $k$, the vast majority of lines rise steeply. In particular, in 12 out of 15 subreddits the outlier users exhibit large increases in toxicity. The subreddits r/imgoingtohellforthis2 and r/wojak are those featuring the largest toxicity contributions by outlier users. On the contrary, outliers in r/hatecrimehoaxes and r/oandaexclusiveforum show marked toxicity reductions, while r/chapotraphouse remains overall stationary. In summary, the results presented in this section highlight that the presence of resentful users who became much more toxic after The Great Ban was not localized within a single or a few subreddits. On the contrary, despite the fact that only a minority of users experienced marked toxicity increases, such adverse reactions were pervasive across most of the analyzed subreddits, which is indicative of a systemic phenomenon.
RQ3: Behavioral dynamics of extremely toxic users
Our analyses in RQ2 revealed the widespread presence of extremely toxic users and quantified the extent of the issue across the subreddits. Here we move forward by providing additional information on their behavior in order to draw insights into their possible impact on the platform.
Subreddits participation. Initially, we explore the subreddit participation habits of these users as a function of their change in toxicity. This analysis aims to ascertain whether each of the most toxic users engaged in a few or many of the banned subreddits. Figure 7 suggests that only a small number of users participated in multiple subreddits. Instead, the vast majority of users participated only in a few subreddits, typically ranging between one and three. Moreover, user toxicity changes do not depend on the number of subreddits in which the user participated.
Figure 8: Density scatterplot of the relationship between changes in toxicity and changes in activity, for each user. The inset provides counts of the number of users in each quadrant of the scatterplot. Only 16% of all users simultaneously exhibited increases in toxicity and activity, as shown by the low number of users in Q1.
Activity. Next, we look into the relationship between changes in toxicity and activity. We measure user activity as the number of posted comments. Conjointly studying user activity and toxicity is important since toxic users pose a meaningful threat only when high toxicity coincides with elevated activity, which holds the potential to cause a substantial impact on the platform. This joint analysis also enables the identification of those users who are both highly toxic and remarkably active, allowing to prioritize moderation interventions towards them. Figure 8 presents a density scatterplot of the changes in toxicity and activity for each considered user. The plot area is divided into four quadrants, corresponding to the combinations of increase/decrease in toxicity/activity. As shown, the bulk of the distribution lies close to the plot origin, meaning that the ban caused minor changes in both toxicity and activity. Moreover, those users who experienced marked changes in toxicity only had negligible changes in activity, and vice versa. The inset in Figure 8 shows the number of users laying in each quadrant of the scatterplot. The most frequent reaction to the ban was a reduction in both toxicity and activity (38% users are in Q3). The opposite behavior is instead the least frequent (16% users are in Q1). Overall this analysis suggests that the resentful users might have had a limited impact on the platform due to their low activity.
Figure 9: Density scatterplot of the relationship between changes in toxicity and changes in score, for each user. The score measures the difference between the total number of upvotes and downvotes received by a user. The inset provides counts of the number of users in each quadrant of the scatterplot. Similarly to Figure 8, only 21% of all users simultaneously exhibited increases in toxicity and score.
Engagement. Finally, we examine the relationship between changes in toxicity and obtained engagement. We measure the engagement obtained by each comment as its Reddit score, computed as the difference between the upvotes and downvotes to that comment. Positive scores denote a positive engagement of the community with a comment, while negative scores generally denote disagreement. The engagement of a user is the mean engagement obtained by its comments. This joint analysis of toxicity and engagement is relevant as it allows an understanding of whether post-ban toxic comments are tolerated, and possibly even encouraged, on Reddit. Figure 9 shows the density scatterplot of this relationship. Similarly to Figure 8, we found that the vast majority of users experienced small changes in toxicity and engagement. Again, the most frequent reaction was a simultaneous reduction of both toxicity and obtained engagement (31% users are in Q3) and the least frequent one was a simultaneous increase of both (21% users are in Q1).
Figure 10: Difference between the score obtained by toxic comments within and outside of the 15 banned subreddits, as a function of comment toxicity. Significance levels of the differences are obtained with non-parametric Mann-Whitney tests for unpaired data: *: $p<0.1$; **: $p<0.05$; ***: $p<0.01$.
Lastly, we analyze the relationship between toxicity and engagement also before the ban, by investigating differences between the engagement obtained within vs outside of the 15 banned subreddits. Figure 10 provides the results of this comparison, highlighting that toxic comments received significantly more engagement within the banned subreddits rather than outside of them (i.e., in the rest of Reddit). This result corroborates that of Figure 9 in that it reinforces the idea that toxicity is not encouraged on Reddit, with the exception of a few toxic subreddits.
Discussion
@@
Our results shed light on the complex effects of The Great Ban, a paramount example of deplatforming that involved around 2,000 subreddits. Among our main findings is that 15.6% of the affected users abandoned Reddit after the ban. Those who remained on the platform reduced their toxicity by 6.6% on average. At the same time however, around 5% of all users markedly increased their toxicity. The presence of such resentful users was widespread across the analyzed subreddits. However, only 16% of users increased both their toxicity and their activity. Similarly, only 21% of users increased their toxicity and also obtained stronger positive feedback from other users. These nuanced results cover new ground on the effects of The Great Ban and, more broadly, on adverse reactions to moderation interventions. Our results also lend themselves to multiple considerations about the design and deployment of effective moderation as well as about the challenges of regulating online platforms.
@@
Effectiveness of the moderation. In literature, the effectiveness of content moderation actions has been primarily assessed in terms of changes that the moderation caused to the activity and toxicity of the affected users (Chandrasekharan et al. 2017; Jhaver et al. 2021). To this regard, our study revealed that a considerable share of toxic users abandoned the platform while the others exhibited a modest reduction in toxicity and a marked reduction in activity. At first glance, these results appear to be indicative of a successful moderation. However, it is necessary to deeply scrutinize these findings in light of potential unintended consequences. For example, the departure of some of the most toxic users from the platform suggests a form of displacement rather than a resolution of the toxicity issue. As recent research pointed out, these users have likely migrated to other online spaces (Horta Ribeiro et al. 2021), raising concerns about the displacement of their toxic interactions instead of their mitigation. These worries are emphasized by the knowledge that users who migrate after facing restrictions on a platform, subsequently engage in even more toxic and aggressive behavior (Horta Ribeiro et al. 2021). In addition, user churn and diminished activity levels post-ban might pose challenges for Reddit, as online platforms thrive on user engagement and interactions for generating revenues (Trujillo and Cresci 2022). Consequently, the apparent success of The Great Ban in mitigating toxicity must be interpreted with caution, considering the potential negative impact on the broader online ecosystem – due to user migrations – and on the platform’s economic model – due to user churn and reduced activity. Future endeavors should aim to strike a balance, devising strategies to curb toxicity without inducing abandonment or substantial decreases in user activity. The quest for effective moderation should align with the overarching goal of cultivating healthier online communities without compromising the safety of other platforms or economic viability, which could otherwise disincentivize platforms to carry out scrupulous moderation.
@@
Divergent reactions to moderation. Our analysis also revealed that, in spite of a modest overall reduction in toxicity, a non-negligible minority of users exhibited large toxicity increases. This result has important implications for the assessment and development of moderation interventions. First, it sheds light on the complexity of user reactions to content moderation, which is a largely underexplored area that requires further investigation (Jhaver et al. 2023). In addition, it surfaces the need for a more nuanced and personalized approach to content moderation. A generic intervention such as The Great Ban – which involved thousands of subreddits and tens of thousands of users – may not effectively address the diverse motivations and behaviors exhibited by the affected users. In our work, this was exemplified by the minority of resentful users who greatly increased their toxicity. Understanding the factors contributing to such divergent responses is paramount for developing effective moderation strategies. Future research and practical applications should delve into user profiling, considering individual characteristics, past behavior, and contextual factors to effectively tailor moderation interventions (Cresci, Trujillo, and Fagni 2022). Moreover, the migration of a subset of toxic users and the widespread presence of users experiencing heightened toxicity raises questions about the potential radicalization effect of moderation and its unintended contribution to the amplification of echo chambers (Horta Ribeiro et al. 2021). It suggests that some users may react negatively to certain moderation actions, possibly leading to more extreme behaviors. These observations highlight the delicate balance required in content moderation, where the aim should not only be that of reducing toxicity locally, but also preventing inadvertent consequences that might exacerbate polarization or radicalization in some user groups.
Impact of resentful users. The finding that the vast majority of resentful users – who exhibited increased toxicity following the ban – did not concurrently escalate their activity offers a novel perspective on the little-studied interplay of toxicity and user activity. Contrary to the suspicion that highly toxic users might also be highly active, our results underscore a limited overlap between these two dimensions of user behavior. This suggests that a considerable portion of users expressing heightened toxicity post-ban did not translate their resentment into increased participation on the platform. In turn, this result partly mitigates the previous concerns about the divergent and adverse reactions to The Great Ban. Furthermore, the observation that users who intensified their toxic behaviors did not receive heightened positive social feedback from their peers resonates with some recent research on the drivers of online hateful and toxic speech. The latter posits that online toxicity may be driven by the pursuit of social approval rather than by a desire to harm (Jiang et al. 2023). To this end, our results demonstrate that the Reddit community did not endorse or support the increased toxicity exhibited by these users. According to the aforementioned theories, toxic users who do not receive the expected levels of social approval should decrease their toxicity over time (Jiang et al. 2023). However, we measured overall stationary toxicity trends over the course of the 7 months after the ban. Our results thus apparently challenge the social approval theory, in support of competing theories according to which at least a subset of toxic and hateful users seek amusement by harassing others, as in trolling behavior (Chandrasekharan et al. 2022). On the one hand, these results call for additional research to identify the types and characteristics of toxic users. On the other hand, they reinforce the need for nuanced and diversified moderation strategies that are capable of addressing the complexity and multiplicity of online user behavior.
@@
Limitations. Our study is based on a large historical dataset of Reddit comments shared by 17K users during a long observation window of 14 months centered around The Great Ban. Due to the nature of our dataset, our findings may be specific to Reddit and the context of The Great Ban. Therefore, caution is needed in generalizing the results to other online platforms or different moderation interventions. Similarly, online platforms are dynamic environments that are subject to continuous changes in user behavior, community norms, and platform policies. Our study covers a specific snapshot in time, albeit relatively long, which nonetheless limits the possibility to carry over our results to different time periods and, partially, also to account for long-term or evolving trends. Another limitation of our work lies in its observational nature, which hinders the possibility of accounting for external events or changes in the broader online ecosystem that may have influenced user behavior independently of The Great Ban. For this reason, care is needed in establishing causal relationships from the findings presented herein. Finally, our dataset lacks comprehensive information about user demographics, motivations, or contextual factors. Accounting for these aspects could provide a more nuanced and actionable interpretation of our results. In this regard, future research could investigate the unexplored interplay between user characteristics and the outcomes of moderation, also as a preliminary step towards the development of targeted and personalized moderation interventions.
@@
Ethical considerations. This research contributes to a deeper understanding of the impact of content moderation, shedding new light on the complexities of user reactions to moderation interventions. This knowledge can inform future developments of effective and nuanced moderation strategies aimed at curbing online toxicity while minimizing unintended consequences. To this end, our work draws attention on the trade off between common versus minority good. In our work this is exemplified by the ethical dilemma faced by moderators who must decide whether to enforce interventions that could possibly harm a minority of users by making them resentful and vindictive, in order to provide a modest benefit to the broader community.
Conclusions

### v5

@@ 4. Analyses and Results
Table 2. Subreddit-wise median toxicity scores for users who abandoned Reddit after the ban (ABA) and for those who remained. For the latter, toxicity scores are computed both before (BEF) and after (AFT) the ban. The ABA vs BEF and BEF vs AFT columns show the effect sizes and the statistical significance levels of the differences in toxicity. The ABA vs BEF column shows that users who abandoned the platform were more toxic than those who remained. The BEF vs AFT column shows that users who remained active experienced a modest toxicity reduction after the ban.
4.1. RQ1: Effectiveness of The Great Ban
@@
Abandoning users. Table 1 highlights a difference of 2,577 users (15.6%) between the OUT-BEFORE and OUT-AFTER datasets, corresponding to users who became inactive on Reddit after The Great Ban. Given that such users were consistently active before the ban, but did not post a single comment in the 7 months following it, we conclude that they abandoned Reddit, possibly migrating to other platforms (Horta Ribeiro et al., 2021). This result thus highlights a first straightforward effect of The Great Ban. In order to draw more insights into this finding, we compare the pre-ban subreddit-wise toxicity of the abandoning users with that of the users who remained active on the platform. Table 2 shows that in 14 out of 15 subreddits, users who later abandoned the platform were more toxic than those who remained after the ban. In other words, toxic users were more likely to abandon the platform after the intervention than less toxic ones. For 12 subreddits the difference in toxicity between abandoning and remaining users is statistically significant ($p<0.05$), according to a non-parametric Mann-Whitney test for unpaired data. Moreover, in 12 subreddits, abandoning users have larger mean absolute deviation (MAD) scores than remaining users, indicating greater variability in toxicity among the former user group.
@@
Remaining users. In addition to causing some users to abandon the platform, the ban might also have caused toxicity changes in those users who remained. Table 2 reports subreddit-wise toxicity scores for the matched set of remaining users, before and after the ban. As shown, the remaining users were on average less toxic after the ban. Specifically, when aggregating results for each subreddit, users from all 15 subreddits slightly decreased their toxicity. The decrease is statistically significant for 11 subreddits ($p<0.1$), according to a non-parametric Wilcoxon test for paired data. Notably, effect sizes for the comparison between remaining users before vs after the ban are smaller than those of the comparison between abandoning vs remaining users. In fact, after The Great Ban, the overall toxicity decreased by 6.57% on average —a modest amount. Moreover, for 13 out of 15 subreddits, toxicity MAD values are larger after the ban than before. This result suggests that the ban increased the degree of variability in the behavior of the remaining users, which we further investigate in the following.
@@
User-level effects. So far, we provided platform- and community-level results, finding that The Great Ban caused 15.6% of core users to abandon the platform. Furthermore, those who remained exhibited a modest average toxicity reduction of 6.57%. To provide a thorough assessment of the intervention, here we also investigate user-level effects by computing the toxicity changes experienced by each of the 13,963 users who remained active on Reddit after the ban. The central panel of Figure 3 presents a slope chart of all user-level toxicity changes, independently of the subreddit in which users participated. Each line corresponds to a single user. Line slopes encode the amount of toxicity reduction or increment. Rising lines are colored with different shades of red depending on their slope, and denote users who increased their toxicity after the ban. Contrarily, decreasing lines are blue-colored and denote users who decreased their toxicity. Figure 3 also includes marginal boxplot toxicity distributions for remaining users before (left-hand side of the slope chart) and after (right-hand side) the ban. The leftmost boxplot presents the toxicity distribution for users who abandoned the platform. Finally, the bottom panel shows the distribution of user-level toxicity changes as a beeswarm plot. This latter plot is useful for highlighting outliers and for studying the two tails of the distribution —i.e., those related to marked toxicity increases (red dots, right-hand side of the beeswarm plot) and decreases (blue dots, left-hand side).
@@
The three boxplots of Figure 3 confirm the overall toxicity trends observed in Table 2. Abandoning users have the largest median toxicity, followed by remaining users before the ban. Out of the three user groups, the remaining users after the ban have the lowest median toxicity. The slope chart in Figure 3 depicts a majority of lines with relatively small positive or negative slopes. These correspond to users who exhibited small toxicity changes —either increases or decreases— after the ban. At the same time, however, the slope chart also features a remarkable number of steep lines, which are related to users who exhibited large toxicity changes. Specifically, we note an overwhelming majority of red-colored steep lines. This implies that, among users who exhibited large toxicity changes, the vast majority increased their toxicity. This result qualitatively describes an undesired effect of The Great Ban, which caused a non-negligible minority of users to become resentful, thus exhibiting much more toxic behaviors. The points where the lines of the slope chart intersect the y axis on the right-hand side of the plot represent the toxicity of the users after the ban. The distribution of the post-ban toxicity scores is also depicted in the right-hand side boxplot. As shown in Figure 3 and as anticipated from the MAD values in Table 2, there is more variability in user toxicity after the ban. Figure 3 clarifies that this increased variability is caused by the increased toxicity of the resentful users.
Figure 3. User-level toxicity change after The Great Ban, for each active user. The slope chart in the central panel reveals a majority of red-colored rising lines, corresponding to a large number of users who drastically increased their toxicity. The beeswarm plot in the bottom panel confirms this effect, presenting more users in the right red-colored tail of the distribution than in the left blue-colored one. Boxplots present marginal distributions for abandoning and remaining user, before and after the intervention.
@@
4.2. RQ2: Extreme user reactions to The Great Ban
@@
Figure 4. User-level toxicity changes for participants in r/the_donald (top) and r/consumeproduct (bottom).
Figure 5. Subreddit-wise changes in toxicity obtained by summing all individual decreasing and increasing contributions, for all users (left panel) and outlier users (right panel). Outlier users are those whose change in toxicity exceeds the threshold $k=0.25$. Based on this criterion, we found no outliers for r/hatecrimehoaxes, which is thus omitted from the right panel figure.
@@
Figure 6. Contribution of the outlier users to the increase/decrease of toxicity in each subreddit. Outlier users are defined as those whose individual change in toxicity exceeds the threshold $k$. As shown, in 12 out of 15 subreddits the outlier users caused large toxicity increases.
@@
In order to specifically investigate the behavior of outlier users —those who experienced marked changes in toxicity— we recompute the previous summations by only considering users whose change in toxicity exceeds a given threshold $k$: $|\Delta t(i)|>k$. This allows us to focus on those users who experienced extreme behavioral changes, be them increases or decreases in toxicity. For example, the right panel of Figure 5 repeats the previous analysis by applying the threshold $k=0.25$. As shown, we obtain strikingly different results. When only considering outlier users, in 13 out of 15 subreddits the contributions of those who increased or decreased their toxicity to the overall subreddit toxicity are very imbalanced. The only exceptions are r/gendercritical who presents relatively balanced contributions (55% vs 45%) and r/hatecrimehoaxes for which no user experienced a change in toxicity exceeding the threshold $k=0.25$. Moreover, in 11 subreddits toxicity increases greatly outweigh decreases. To generalize these findings, we repeated the analysis by varying the value of $k$ from 0 to 1, with a step of 0.01. Results are shown in Figure 6 where each line corresponds to a subreddit and reports the % difference $\Delta t^{+}-\Delta t^{-}$ between the contributions of the two tails, for each value of $k$. In the figure, only a few notable subreddits are highlighted, while the rest are grey-colored so as to reveal the overall trend. As shown, even for small values of the threshold $k$, the vast majority of lines rise steeply. In particular, in 12 out of 15 subreddits the outlier users exhibit large increases in toxicity. The subreddits r/imgoingtohellforthis2 and r/wojak are those featuring the largest toxicity contributions by outlier users. On the contrary, outliers in r/hatecrimehoaxes and r/oandaexclusiveforum show marked toxicity reductions, while r/chapotraphouse remains overall stationary. However, the spike observed in r/imgoingtohellforthis2, as well as the decline of r/oandaexclusiveforum, may be attributed to the relatively small number of users (less than 100). In summary, the results presented in this section highlight that the presence of resentful users who became much more toxic after The Great Ban was not localized within a single or a few subreddits. On the contrary, despite the fact that only a minority of users experienced marked toxicity increases, such adverse reactions were pervasive across most of the analyzed subreddits, which is indicative of a systemic phenomenon.
5. Discussion
@@
Our results shed light on the complex effects of The Great Ban, a paramount example of deplatforming that involved around 2,000 subreddits. Among our main findings is that 15.6% of the affected users abandoned Reddit after the ban. Those who remained on the platform reduced their toxicity by 6.6% on average. At the same time however, around 5% of all users markedly increased their toxicity. The presence of such resentful users was widespread across the analyzed subreddits. These nuanced results cover new ground on the effects of The Great Ban and, more broadly, on adverse reactions to moderation interventions. Our results also lend themselves to multiple considerations about the design and deployment of effective moderation as well as about the challenges of regulating online platforms.
@@
Effectiveness of the moderation. In literature, the effectiveness of content moderation actions has been primarily assessed in terms of changes that the moderation caused to the activity and toxicity of the affected users (Chandrasekharan et al., 2017; Jhaver et al., 2021). To this regard, our study revealed that a considerable share of toxic users abandoned the platform while the others exhibited a modest reduction in toxicity and a marked reduction in activity. At first glance, these results appear to be indicative of a successful moderation. However, it is necessary to deeply scrutinize these findings in light of potential unintended consequences. For example, the departure of some of the most toxic users from the platform suggests a form of displacement rather than a resolution of the toxicity issue. As recent research pointed out, these users have likely migrated to other online spaces (Horta Ribeiro et al., 2021), raising concerns about the displacement of their toxic interactions instead of their mitigation. These worries are emphasized by the knowledge that users who migrate after facing restrictions on a platform, subsequently engage in even more toxic and aggressive behavior (Horta Ribeiro et al., 2021). In addition, user churn and diminished activity levels post-ban might pose challenges for Reddit, as online platforms thrive on user engagement and interactions for generating revenues (Trujillo and Cresci, 2022). Consequently, the apparent success of The Great Ban in mitigating toxicity must be interpreted with caution, considering the potential negative impact on the broader online ecosystem —due to user migrations— and on the platform’s economic model —due to user churn and reduced activity. Future endeavors should aim to strike a balance, devising strategies to curb toxicity without inducing abandonment or substantial decreases in user activity. The quest for effective moderation should align with the overarching goal of cultivating healthier online communities without compromising the safety of other platforms or economic viability, which could otherwise disincentivize platforms to carry out scrupulous moderation.
@@
Divergent reactions to moderation. Our analysis also revealed that, in spite of a modest overall reduction in toxicity, a non-negligible minority of users exhibited large toxicity increases. This result has important implications for the assessment and development of moderation interventions. First, it sheds light on the complexity of user reactions to content moderation, which is a largely underexplored area that requires further investigation (Jhaver et al., 2023). In addition, it surfaces the need for a more nuanced and personalized approach to content moderation. A generic intervention such as The Great Ban —which involved thousands of subreddits and tens of thousands of users— may not effectively address the diverse motivations and behaviors exhibited by the affected users. In our work, this was exemplified by the minority of resentful users who greatly increased their toxicity. Understanding the factors contributing to such divergent responses is paramount for developing effective moderation strategies. Future research and practical applications should delve into user profiling, considering individual characteristics, past behavior, and contextual factors to effectively tailor moderation interventions (Cresci et al., 2022). Moreover, the migration of a subset of toxic users and the widespread presence of users experiencing heightened toxicity raises questions about the potential radicalization effect of moderation and its unintended contribution to the amplification of echo chambers (Horta Ribeiro et al., 2021). It suggests that some users may react negatively to certain moderation actions, possibly leading to more extreme behaviors. These observations highlight the delicate balance required in content moderation, where the aim should not only be that of reducing toxicity locally, but also preventing inadvertent consequences that might exacerbate polarization or radicalization in some user groups.
@@
Limitations. Our study is based on a large historical dataset of Reddit comments shared by 17K users during a long observation window of 14 months centered around The Great Ban. Due to the nature of our dataset, our findings may be specific to Reddit and the context of The Great Ban. Therefore, caution is needed in generalizing the results to other online platforms or different moderation interventions. Similarly, online platforms are dynamic environments that are subject to continuous changes in user behavior, community norms, and platform policies. Our study covers a specific snapshot in time, albeit relatively long, which nonetheless limits the possibility to carry over our results to different time periods and, partially, also to account for long-term or evolving trends. Moreover, toxicity scores of comments were obtained via the machine learning model Detoxify, which in itself might have introduced a biased evaluation of toxicity. Another limitation of our work lies in its observational nature, which hinders the possibility of accounting for external events or changes in the broader online ecosystem that may have influenced user behavior independently of The Great Ban. For this reason, care is needed in establishing causal relationships from the findings presented herein. Future work could adopt more sophisticated causal inference techniques, for instance difference-in-differences or interrupted time series. However, finding suitable control subreddits and taking into account exogenous events remains a significant challenge (Trujillo and Cresci, 2022). Finally, our dataset lacks comprehensive information about user demographics, motivations, or contextual factors. Accounting for these aspects could provide a more nuanced and actionable interpretation of our results. In this regard, future research could investigate the unexplored interplay between user characteristics and the outcomes of moderation, also as a preliminary step towards the development of targeted and personalized moderation interventions (Cresci et al., 2022).
@@
Ethical considerations. This research contributes to a deeper understanding of the impact of content moderation, shedding new light on the complexities of user reactions to moderation interventions. This knowledge can inform future developments of effective and nuanced moderation strategies aimed at curbing online toxicity while minimizing unintended consequences. To this end, our work draws attention on the trade off between common versus minority good. In our work this is exemplified by the ethical dilemma faced by moderators who must decide whether to enforce interventions that could possibly harm a minority of users by making them resentful, in order to provide a modest benefit to the broader community.
6. Conclusions